# Acylindrical hyperbolicity and the centers of Artin groups that are not free of infinity

### v2

@@ 3. Reduced clique-cube complexes and actions on them
We show that $C_{\Gamma,\Gamma_{0}}$ is convex in $C_{\Gamma}$. Let $gA_{U}$ be a vertex of $C_{\Gamma,\Gamma_{0}}$. Let $L$ be the link of $gA_{U}$ in $C_{\Gamma}$ and let $L^{\prime}$ be the link of $gA_{U}$ in $C_{\Gamma,\Gamma_{0}}$. According to Theorem 3.5, it is sufficient to show that $L^{\prime}$ is a full subcomplex of $L$. Let $\Delta$ be a $d$-dimensional simplex of $L$ whose vertices are in $L^{\prime}$. Let $C_{\Delta}$ be a $(d+1)$-dimensional cube of $C_{\Gamma}$ which corresponds to $\Delta$. Let $gA_{U_{1}},\ldots,gA_{U_{d+1}}$ be vertices of $C_{\Delta}$ that are connected to $gA_{U}$ by a $1$-dimensional cube of $C_{\Gamma}$. Note that $C_{\Delta}=[gA_{\cap_{i=1}^{d+1}U_{i}},gA_{\cup_{i=1}^{d+1}U_{i}}]$. Since $gA_{U_{i}}\in C_{\Gamma,\Gamma_{0}}$, $U_{i}\supset V(\Gamma_{0})$ for every $i$. Therefore, $\cap_{i=1}^{d+1}U_{i}\supset V(\Gamma_{0})$. Since $C_{\Delta}$ is contained in $C_{\Gamma,\Gamma_{0}}$, $\Delta$ is contained in $L^{\prime}$. It follows that $L^{\prime}$ is a full subcomplex of $L$.
Moreover, $C_{\Gamma,\Gamma_{0}}$ is connected (Lemma 3.2), thus it is CAT(0) as a subspace of a CAT(0) space $C_{\Gamma}$. Completeness follows from the fact that $C_{\Gamma,\Gamma_{0}}$ is a finite dimensional cube complex ([7, Theorem 7.19]). ∎
@@
When $\Gamma$ is a clique, $C_{\Gamma,\Gamma_{0}}=\{A_{\Gamma_{0}}\}$. In this case, the lemma is trivial. We consider the case where $\Gamma$ is not a clique. We fix a join decomposition
@@
where $\Gamma_{0}$ is the maximal clique factor of $\Gamma$ and $\Gamma_{i}$ are indecomposable factors with at least two vertices for $i\geq 1$. We note that $\Gamma_{1}\ast\cdots\ast\Gamma_{k}\neq\emptyset$ since $\Gamma$ is not a clique.
Let $x\in C_{\Gamma,\Gamma_{0}}$. Let $C$ be a cube of $C_{\Gamma,\Gamma_{0}}$ such that $x\in C$. Without loss of generality, we set $C=[A_{U},A_{T}]$ where $U\subset T$, $U,T\in\Delta_{\Gamma,\Gamma_{0}}$. By induction on $|T\setminus U|$, we show that $\mathrm{Hull}(orb(x))=C_{\Gamma,\Gamma_{0}}$.
When $|T\setminus U|=0$, then $x=A_{U}$. It is sufficient to show that $\mathrm{Hull}(orb(A_{U}))=\mathrm{Hull}(orb(A_{V(\Gamma_{0})}))$. Indeed, if $\mathrm{Hull}(orb(A_{U}))=\mathrm{Hull}(orb(A_{V(\Gamma_{0})}))$ for every $U\in\Delta_{\Gamma,\Gamma_{0}}$, then $\mathrm{Hull}(orb(A_{V(\Gamma_{0})}))$ contains every vertex of $C_{\Gamma,\Gamma_{0}}$ and thus equals to $C_{\Gamma,\Gamma_{0}}$.
First, we show that $\mathrm{Hull}(orb(A_{U}))\supset\mathrm{Hull}(orb(A_{V(\Gamma_{0})}))$. If $V(\Gamma_{0})\neq U$, we take $u\in U\setminus V(\Gamma_{0})$. There exists $i\in\{1,\ldots,k\}$ such that $u\in\Gamma_{i}$. Since $\Gamma_{i}$ is indecomposable as $\Gamma_{i}=u\ast\Gamma^{\prime}_{i}$, there exists a vertex $s\neq u$ of $\Gamma_{i}$ such that $s$ and $u$ are not connected by an edge in $\Gamma_{i}$. Let
$U^{\prime}=\{v\in V_{\Gamma}\mid\text{ $v$ and $s$ are connected by an edge in$\Gamma$}\}\cap U.$
We note that ${U^{\prime}}$ does not contain $u$ and $s$, and that ${U^{\prime}}\subsetneq U$. We also note that ${U^{\prime}}\cup\{s\}$ spans a clique, but $U\cup\{s\}$ does not. Then,
$A_{U},\ A_{U^{\prime}},\ A_{{U^{\prime}}\cup\{s\}},\ sA_{U^{\prime}},\ sA_{U}$
is a local geodesic, and thus a geodesic, since $C_{\Gamma,\Gamma_{0}}$ is CAT(0) (Lemma 3.6). This geodesic is contained in $\mathrm{Hull}(orb(A_{U}))$, since two endpoints are in $\mathrm{Hull}(orb(A_{U}))$. Therefore, $A_{U^{\prime}}\in\mathrm{Hull}(orb(A_{U}))$, and thus $\mathrm{Hull}(orb(A_{U^{\prime}}))\subset\mathrm{Hull}(orb(A_{U}))$. If $U^{\prime}=V(\Gamma_{0})$, then we have nothing to do. If ${U^{\prime}}\supsetneq V(\Gamma_{0})$, we repeat the same argument until we get $\mathrm{Hull}(orb(A_{U}))\supset\mathrm{Hull}(orb(A_{V(\Gamma_{0})}))$.
Next, we show that $\mathrm{Hull}(orb(A_{U}))\subset\mathrm{Hull}(orb(A_{V(\Gamma_{0})}))$. If $V(\Gamma_{0})\neq U$, we take $u\in U\setminus V(\Gamma_{0})$. Then
@@
is a geodesic. It follows that $\mathrm{Hull}(orb(A_{U}))\subset\mathrm{Hull}(orb(A_{U\setminus\{u\}}))$. If $U\setminus\{u\}=V(\Gamma_{0})$, then the proof is finished. If $U\setminus\{u\}\supsetneq V(\Gamma_{0})$, we repeat the same argument until we get $\mathrm{Hull}(orb(A_{U}))\subset\mathrm{Hull}(orb(A_{V(\Gamma_{0})}))$.
When $|T\setminus U|>0$, there exists $s\in T\setminus U$. Then $C\cap sC$ contains a face spanned by $A_{s}$ and $A_{T}$. The geodesic from $x$ to $sx$ passes through this face. Let $y$ be the intersection point. Then $y\in\mathrm{Hull}(orb(x))$, and thus $\mathrm{Hull}(orb(y))\subset\mathrm{Hull}(orb(x))$. By induction, $\mathrm{Hull}(orb(y))=C_{\Gamma,\Gamma_{0}}$. ∎
@@
Suppose that $\Gamma$ is not a clique. Then, a finite normal subgroup of $A_{\Gamma}$ is in $A_{\Gamma_{0}}$. In particular, any finite subgroup of the center of $A_{\Gamma}$ is in $A_{\Gamma_{0}}$. Also, if $A_{\Gamma}$ is isomorphic to a direct product $A_{1}\times A_{2}$ and $A_{1}$ is finite, then $A_{1}$ is in $A_{\Gamma_{0}}$ and satisfies $A_{\Gamma_{0}}=A_{1}\times(A_{\Gamma_{0}}\cap A_{2})$.
@@
Because $N$ is finite and $C_{\Gamma,\Gamma_{0}}$ is a complete CAT(0) space, we have $Fix(N)\neq\emptyset$. Take any $x\in Fix(N)$. Then, $A_{\Gamma}x\subset Fix(N)$. Indeed, the normality of $N$ implies that, for any $g\in A_{\Gamma}$ and $n\in N$, there exists $n^{\prime}\in N$ such that $ng=gn^{\prime}$. Therefore, we have $ngx=gn^{\prime}x=gx$. Because $C_{\Gamma,\Gamma_{0}}$ is CAT(0), $Fix(N)$ is convex. Hence, we have $Hull(A_{\Gamma}x)\subset Fix(N)$. By Lemma 3.7, we have $Hull(A_{\Gamma}x)=C_{\Gamma,\Gamma_{0}}$. Hence, $Fix(N)=C_{\Gamma,\Gamma_{0}}$. In particular, $A_{\Gamma_{0}}\in Fix(N)$. Hence, $N$ must be in $A_{\Gamma_{0}}$. ∎
Finally, we state a lemma related to the local geometry of reduced clique-cube complexes. The following is a nontrivial variant for $C_{\Gamma,\Gamma_{0}}$ of [29, Lemma 5.2] for $C_{\Gamma}$.
@@
Let $s,t\in V(\Gamma_{\ast})$ belong to mutually different indecomposable factors of $\Gamma_{\ast}$. Suppose that we have $d\in\mathbb{N}$ and mutually different $w_{1},\ldots,w_{d-1}\in V(\Gamma_{0})$ such that $(w_{i-1},w_{i})\in E(\Gamma)$ with label greater than $2$ for any $i\in\{1,\ldots,d\}$ where we set $w_{0}:=s$ and $w_{d}:=t$. Set $\tau:=w_{0}w_{1}\cdots w_{d}\in A_{\Gamma}$. Let $U\in\Delta_{\Gamma,\Gamma_{0}}$ with $s,t\in U$. Then, there exists no square in $C_{\Gamma,\Gamma_{0}}$ containing both edges $[A_{U},A_{U\setminus\{s\}}]$ and $[A_{U},\tau A_{U\setminus\{t\}}]$, that is, there exists no $g\in A_{U\setminus\{s\}}$ such that $gA_{U\setminus\{t\}}=\tau A_{U\setminus\{t\}}$.
@@
Figure 3. Squares around $A_{U}$.
@@
Denote by $\bar{\tau}$ the image of $\tau$ by the natural projection $A_{\Gamma}\to W_{\Gamma}$. Since $w_{0},w_{1},\ldots,w_{d}$ are mutually different, the word $w_{0}w_{1}\cdots w_{d}$ is a reduced expression of $\bar{\tau}$. Moreover this is a unique word representing $\bar{\tau}$ because $\bar{\tau}$ is regarded as a Coxeter element in $W_{\{w_{0},w_{1},\ldots,w_{d}\}}$, $w_{i-1}$ and $w_{i}$ cannot commute for any $i\in\{1,\ldots,d\}$ (see for example [39]).
Assume that we have $g\in A_{U\setminus\{s\}}$ such that $gA_{U\setminus\{t\}}=\tau A_{U\setminus\{t\}}$. Then we have $\tau\in A_{U\setminus\{s\}}A_{U\setminus\{t\}}$ and thus $\bar{\tau}\in W_{U\setminus\{s\}}W_{U\setminus\{t\}}$. Take $a_{1},\ldots,a_{d_{1}}\in U\setminus\{s\}$ and $b_{1},\ldots,b_{d_{2}}\in U\setminus\{t\}$ such that the word $a_{1}\cdots a_{d_{1}}b_{1}\cdots b_{d_{2}}$ represents $\bar{\tau}$. Then by the deletion property for Coxeter groups (see for example [3]), the word $a_{1}\cdots a_{d_{1}}b_{1}\cdots b_{d_{2}}$ contains a reduced expression for $\tau$ as a subword. In such any subword, the letter $t$ precedes $s$. This contradicts the fact that $w_{0}=s$ precedes $w_{d}=t$ in the unique reduced expresssion $w_{0}w_{1}\cdots w_{d}$ of $\bar{\tau}$. ∎
@@
Let $A_{\Gamma}$ be an Artin group associated with a defining graph $\Gamma$ with the vertex set $V(\Gamma)=V$ and the edge set $E(\Gamma)=E$, as in Section 2.2. In this section, we suppose that $\Gamma$ is not a clique. We show our main theorem (Theorem 1.2), that is, the following are equivalent:
@@
    $A_{\Gamma}$ is irreducible, that is, $\Gamma$ cannot be decomposed as a join of two subgraphs such that all edges between them are labeled by $2$;
@@
    $A_{\Gamma}$ is virtually directly indecomposable, that is, it cannot be decomposed as a direct product of two infinite subgroups.
@@
We give a proof of $(1)\Rightarrow(2)$ in Theorem 1.2. Suppose that $A_{\Gamma}$ is irreducible, that is, $\Gamma^{t}$ is connected. We will construct an element $\gamma\in A_{\Gamma}$ and show that $\gamma$ is a WPD contracting element with respect to the action on the reduced clique-cube complex $C_{\Gamma,\Gamma_{0}}$. When $\Gamma_{0}$ is empty, a WPD contracting element is given in [29]. Our $\gamma$ in this paper is constructed more delicately than in [29].
We consider the unique join decomposition $\Gamma=\Gamma_{0}\ast\Gamma_{\ast}=\Gamma_{0}\ast(\ast_{1\leq i\leq k}\Gamma_{i})$ where $\Gamma_{0}$ is the maximal clique factor and $\Gamma_{i}$ ($1\leq i\leq k$) is an indecomposable factor with at least two vertices (see Definition 2.18). We set $V_{i}=V(\Gamma_{i})$ and $E_{i}=E(\Gamma_{i})$ for each $i\in\{0,1,\ldots,k\}$. Then, for every $i\neq 0$, the complement graph $(\Gamma_{i})^{c}$ of $\Gamma_{i}$ is connected and the complement graph $(\Gamma_{\ast})^{c}$ of $\Gamma_{\ast}$ is the disjoint union of connected components as $(\Gamma_{\ast})^{c}=\bigsqcup_{1\leq i\leq k}(\Gamma_{i})^{c}$ by Lemma 2.17.
@@
  • the vertex set $V(Q(\Gamma,\Gamma_{0}))=\{V_{i}\}_{1\leq i\leq k}$, and
  • the edge set $E(Q(\Gamma,\Gamma_{0}))=\{(V_{i},V_{j}),(V_{j},V_{i})\mid 1\leq i<j\leq k,\\ (\Gamma_{i})^{t}\text{ and }(\Gamma_{j})^{t}\text{ are connected in }(\Gamma_{0}\ast\Gamma_{i}\ast\Gamma_{j})^{t}\}$.
@@
  • there exists a path consisting of edges with label greater than $2$ between a vertex of $\Gamma_{i}$ and a vertex of $\Gamma_{j}$ in $\Gamma_{0}\ast\Gamma_{i}\ast\Gamma_{j}$.
@@
For each $i,j$ with $i<j$ and $(V_{i},V_{j})\in E(T)$, take a path $p_{i,j}=(w_{i,j,0},w_{i,j,1},\ldots,w_{i,j,d(i,j)})$ consisting of edges with label greater than $2$ such that $w_{i,j,1},\ldots,w_{i,j,d(i,j)-1}\in\Gamma_{0}$ if $d(i,j)\geq 2$, $w_{i,j,0}\in\Gamma_{i}$ and $w_{i,j,d(i,j)}\in\Gamma_{j}$, and set $p_{j,i}=(w_{j,i,0},w_{j,i,1},\ldots,w_{j,i,d(j,i)}):=(w_{i,j,d(i,j)},\ldots,w_{i,j,1},w_{i,j,0})$. For any $i,j\in\{1,\ldots,k\}$ with $(V_{i},V_{j})\in E(T)$, set $s_{i,j}:=w_{i,j,0}$ in $\Gamma_{i}$, $t_{i,j}:=w_{i,j,d(i,j)}$ in $\Gamma_{j}$ and $\tau_{i,j}=w_{i,j,0}w_{i,j,1}\cdots w_{i,j,d(i,j)}$ in $A_{\Gamma}$, thus $s_{j,i}=t_{i,j}$ in $\Gamma_{j}$, $t_{j,i}=s_{i,j}$ in $\Gamma_{i}$ and $\tau_{j,i}=w_{i,j,d(i,j)}\cdots w_{i,j,1}w_{i,j,0}$ in $A_{\Gamma}$.
The following is almost same with [29, Lemma 6.1].

### v3

@@ 3. Reduced clique-cube complexes and actions on them
We show that $C_{\Gamma,\Gamma_{0}}$ is convex in $C_{\Gamma}$. Let $gA_{U}$ be a vertex of $C_{\Gamma,\Gamma_{0}}$. Let $L$ be the link of $gA_{U}$ in $C_{\Gamma}$ and let $L^{\prime}$ be the link of $gA_{U}$ in $C_{\Gamma,\Gamma_{0}}$. According to Theorem 3.5 and Lemma 3.2, it is sufficient to show that $L^{\prime}$ is a full subcomplex of $L$. Let $\Delta$ be a $d$-dimensional simplex of $L$ whose vertices are in $L^{\prime}$. Let $C_{\Delta}$ be the $(d+1)$-dimensional cube of $C_{\Gamma}$ corresponding to $\Delta$. Let $gA_{U_{1}},\ldots,gA_{U_{d+1}}$ be vertices of $C_{\Delta}$ that are connected to $gA_{U}$ by a $1$-dimensional cube of $C_{\Gamma}$. Note that $C_{\Delta}=[gA_{\cap_{i=1}^{d+1}U_{i}},gA_{\cup_{i=1}^{d+1}U_{i}}]$. Since $gA_{U_{i}}\in C_{\Gamma,\Gamma_{0}}$, we have $U_{i}\supset V_{0}$ for every $i$. Therefore, $\cap_{i=1}^{d+1}U_{i}\supset V_{0}$. Since $C_{\Delta}$ is contained in $C_{\Gamma,\Gamma_{0}}$, $\Delta$ is contained in $L^{\prime}$. It follows that $L^{\prime}$ is a full subcomplex of $L$.
The other assertions follow from the convexity of $C_{\Gamma,\Gamma_{0}}$ in $C_{\Gamma}$ and Theorem 2.13. ∎
@@
When $\Gamma$ is a clique, $C_{\Gamma,\Gamma_{0}}=\{A_{V_{0}}\}$. In this case, the lemma is trivial. We now treat the case where $\Gamma$ is not a clique. We fix a join decomposition
@@
where $\Gamma_{0}$ is the maximal clique factor of $\Gamma$ and $\Gamma_{i}$ are indecomposable factors with at least two vertices for $i\geq 1$. We note that $\Gamma_{1}\ast\cdots\ast\Gamma_{k}\neq\emptyset$, since $\Gamma$ is not a clique.
Let $x\in C_{\Gamma,\Gamma_{0}}$. Let $C$ be a cube of $C_{\Gamma,\Gamma_{0}}$ with $x\in C$. Without loss of generality, we set $C=[A_{U},A_{T}]$ where $U\subset T$, $U,T\in\Delta_{\Gamma,\Gamma_{0}}$. By induction on #$\#(T\setminus U)$, we show that $\mathrm{Hull}(orb(x))=C_{\Gamma,\Gamma_{0}}$.
When #$\#(T\setminus U)=0$, then $x=A_{U}$. It is sufficient to show that $\mathrm{Hull}(orb(A_{U}))=\mathrm{Hull}(orb(A_{V_{0}}))$. Indeed, if $\mathrm{Hull}(orb(A_{U}))=\mathrm{Hull}(orb(A_{V_{0}}))$ for every $U\in\Delta_{\Gamma,\Gamma_{0}}$, then $\mathrm{Hull}(orb(A_{V_{0}}))$ contains every vertex of $C_{\Gamma,\Gamma_{0}}$ and thus equals $C_{\Gamma,\Gamma_{0}}$.
First, we show that $\mathrm{Hull}(orb(A_{U}))\supset\mathrm{Hull}(orb(A_{V_{0}}))$. If $V_{0}\neq U$, we take $u\in U\setminus V_{0}$. There exists $i\in\{1,\ldots,k\}$ such that $u\in V(\Gamma_{i})$. Since $\Gamma_{i}$ is indecomposable, we have a vertex $s\neq u$ of $\Gamma_{i}$ such that $s$ and $u$ are not connected by an edge in $\Gamma_{i}$. Note that $s\notin U$. Let
$U^{\prime}=\{v\in U\mid\text{ $v$ and $s$ are connected by an edge in $\Gamma$}\}.$
We note that ${U^{\prime}}$ does not contain $u$. Hence we have $V_{0}\subset{U^{\prime}}\subsetneq U$. We also note that ${U^{\prime}}\sqcup\{s\}$ spans a clique but $U\sqcup\{s\}$ does not. Then, the sequence of vertices
$A_{U},\ A_{U^{\prime}},\ A_{{U^{\prime}}\sqcup\{s\}},\ sA_{U^{\prime}},\ sA_{U}$
forms a local geodesic, which is a geodesic, since $C_{\Gamma,\Gamma_{0}}$ is CAT(0) (Lemma 3.6). This geodesic is contained in $\mathrm{Hull}(orb(A_{U}))$, since the two endpoints are in $\mathrm{Hull}(orb(A_{U}))$. Therefore, $A_{U^{\prime}}\in\mathrm{Hull}(orb(A_{U}))$, and thus $\mathrm{Hull}(orb(A_{U^{\prime}}))\subset\mathrm{Hull}(orb(A_{U}))$. If $U^{\prime}=V_{0}$, then we have nothing further to show. If ${U^{\prime}}\supsetneq V_{0}$, we repeat the same argument until we obtain $\mathrm{Hull}(orb(A_{U}))\supset\mathrm{Hull}(orb(A_{V_{0}}))$.
Next, we show that $\mathrm{Hull}(orb(A_{U}))\subset\mathrm{Hull}(orb(A_{V_{0}}))$. If $V_{0}\neq U$, we take $u\in U\setminus V_{0}$. Then
@@
is a geodesic. It follows that $\mathrm{Hull}(orb(A_{U}))\subset\mathrm{Hull}(orb(A_{U\setminus\{u\}}))$. If $U\setminus\{u\}=V_{0}$, then the proof is finished. If $U\setminus\{u\}\supsetneq V_{0}$, we repeat the same argument until we obtain $\mathrm{Hull}(orb(A_{U}))\subset\mathrm{Hull}(orb(A_{V_{0}}))$.
When #$\#(T\setminus U)>0$, there exists $s\in T\setminus U$. Then $C\cap sC$ contains a face spanned by $A_{U\sqcup\{s\}}$ and $A_{T}$. The geodesic from $x$ to $sx$ passes through this face. Let $y$ be the intersection point. Then $y\in\mathrm{Hull}(orb(x))$, and thus $\mathrm{Hull}(orb(y))\subset\mathrm{Hull}(orb(x))$. By induction, $\mathrm{Hull}(orb(y))=C_{\Gamma,\Gamma_{0}}$. ∎
@@
Suppose that $\Gamma$ is not a clique. Then, any finite normal subgroup of $A_{\Gamma}$ is contained in $A_{V_{0}}$. In particular, any finite subgroup of the center of $A_{\Gamma}$ is contained in $A_{V_{0}}$. Also, if $A_{\Gamma}$ is decomposed as a direct product $A_{\Gamma}=A_{1}\times A_{2}$ and $A_{1}$ is finite, then $A_{1}$ is contained in $A_{V_{0}}$ and we have $A_{V_{0}}=A_{1}\times(A_{V_{0}}\cap A_{2})$.
@@
Because $N$ is finite and $C_{\Gamma,\Gamma_{0}}$ is a complete CAT(0) space, we have $Fix(N)\neq\emptyset$. Take any $x\in Fix(N)$. Then, $orb(x)=A_{\Gamma}x\subset Fix(N)$. Indeed, the normality of $N$ implies that, for any $g\in A_{\Gamma}$ and $n\in N$, there exists $n^{\prime}\in N$ such that $ng=gn^{\prime}$. Therefore, we have $ngx=gn^{\prime}x=gx$. Because $C_{\Gamma,\Gamma_{0}}$ is CAT(0), $Fix(N)$ is convex. Hence, we have $Hull(orb(x))\subset Fix(N)$. By Lemma 3.7, we have $Hull(orb(x))=C_{\Gamma,\Gamma_{0}}$. Hence, $Fix(N)=C_{\Gamma,\Gamma_{0}}$. In particular, $A_{V_{0}}\in Fix(N)$. Hence, $N$ must be contained in $A_{V_{0}}$. ∎
Finally, we state a lemma related to the local geometry of reduced clique-cube complexes. The following is a nontrivial variant in the setting of $C_{\Gamma,\Gamma_{0}}$ of [29, Lemma 5.2] concerning $C_{\Gamma}$.
@@
Let $s,t\in V(\Gamma_{\ast})$ belong to mutually different indecomposable factors of $\Gamma_{\ast}$. Suppose that we have $d\in\mathbb{N}$ and mutually different $w_{0},w_{1},\ldots,w_{d}\in V$ such that $w_{0}=s$, $w_{d}=t$ and $(w_{i-1},w_{i})\in E$ with label greater than $2$ for any $i\in\{1,\ldots,d\}$. Moreover, suppose that $w_{i}\in V_{0}$ if $0<i<d$. Set $\tau:=w_{0}w_{1}\cdots w_{d}\in A_{\Gamma}$. Let $U\in\Delta_{\Gamma,\Gamma_{0}}$ with $s,t\in U$. Then, there exists no square in $C_{\Gamma,\Gamma_{0}}$ containing both edges $[A_{U},A_{U\setminus\{s\}}]$ and $[A_{U},\tau A_{U\setminus\{t\}}]$. That is, there exists no $g\in A_{U\setminus\{s\}}$ such that $gA_{U\setminus\{t\}}=\tau A_{U\setminus\{t\}}$.
@@
Figure 6. Squares around $A_{U}$ where $s=w_{0}$, $t=w_{d}$ and $\tau=w_{0}w_{1}\cdots w_{d}$.
@@
Denote by $\bar{\tau}$ the image of $\tau$ by the natural projection $A_{\Gamma}\to W_{\Gamma}$. Since $w_{0},w_{1},\ldots,w_{d}$ are mutually different, the word $w_{0}w_{1}\cdots w_{d}$ is a reduced expression of $\bar{\tau}$. Moreover this is a unique reduced word representing $\bar{\tau}$ because $\bar{\tau}$ is regarded as a Coxeter element in $W_{\{w_{0},w_{1},\ldots,w_{d}\}}$ and two elements $w_{i-1}$ and $w_{i}$ cannot commute for any $i\in\{1,\ldots,d\}$ (see, for example, [39]).
Assume that we have $g\in A_{U\setminus\{s\}}$ such that $gA_{U\setminus\{t\}}=\tau A_{U\setminus\{t\}}$. Then we have $\tau\in A_{U\setminus\{s\}}A_{U\setminus\{t\}}$, and thus $\bar{\tau}\in W_{U\setminus\{s\}}W_{U\setminus\{t\}}$. Take $a_{1},\ldots,a_{d_{1}}\in U\setminus\{s\}$ and $b_{1},\ldots,b_{d_{2}}\in U\setminus\{t\}$ such that the word $a_{1}\cdots a_{d_{1}}b_{1}\cdots b_{d_{2}}$ represents $\bar{\tau}$. Then, by the deletion property for Coxeter groups (see, for example, [3]), the word $a_{1}\cdots a_{d_{1}}b_{1}\cdots b_{d_{2}}$ contains a reduced expression for $\bar{\tau}$ as a subword. In any such subword, the letter $t$ precedes $s$. This contradicts the fact that $w_{0}=s$ precedes $w_{d}=t$ in the unique reduced expression $w_{0}w_{1}\cdots w_{d}$ of $\bar{\tau}$. ∎
@@
We show our main theorem (Theorem 1.2). In fact, we prove the following:
Theorem 4.1.
@@
    $A_{\Gamma}$ is irreducible;
@@
    $A_{\Gamma}$ is virtually directly indecomposable; that is, it cannot be decomposed as a direct product of two infinite subgroups.
In this section, let $A_{\Gamma}$ be an Artin group associated with a defining graph $\Gamma$ with the vertex set $V(\Gamma)=V$ and the edge set $E(\Gamma)=E$, as in Section 2.2. We suppose that $\Gamma$ is not a clique.
@@
We give a proof of $(1)\Rightarrow(2)$ in Theorem 4.1. Suppose that $A_{\Gamma}$ is irreducible; that is, $\Gamma^{t}$ is connected. By noting that $\Gamma$ is not a clique, we consider the unique join decomposition $\Gamma=\Gamma_{0}\ast\Gamma_{\ast}=\Gamma_{0}\ast(\ast_{1\leq i\leq k}\Gamma_{i})$ where $\Gamma_{0}$ is the maximal clique factor and each $\Gamma_{i}$ ($1\leq i\leq k$) is an indecomposable factor with at least two vertices (see Definition 2.18). We set $V_{i}=V(\Gamma_{i})$ and $E_{i}=E(\Gamma_{i})$ for each $i\in\{0,1,\ldots,k\}$. Also, we set $V_{\ast}=V(\Gamma_{\ast})$. Then, for every $i\neq 0$, the complement graph $(\Gamma_{i})^{c}$ of $\Gamma_{i}$ is connected and the complement graph $(\Gamma_{\ast})^{c}$ of $\Gamma_{\ast}$ is the disjoint union of connected components: $(\Gamma_{\ast})^{c}=\bigsqcup_{1\leq i\leq k}(\Gamma_{i})^{c}$, by Lemma 2.17. When $k=1$ and $\Gamma_{0}=\emptyset$, $C_{\Gamma,\Gamma_{0}}=C_{\Gamma}$ and a WPD contracting element is given in [14]. When $k>1$ and $\Gamma_{0}=\emptyset$, a WPD contracting element is given in [29]. Our goal in this section is to give a WPD contracting element with respect to the action on the reduced clique-cube complex $C_{\Gamma,\Gamma_{0}}$ in our general setting. Indeed, we explicitly construct three elements: $\gamma_{\natural}\in A_{V_{\ast}}$, and $\gamma_{\sharp},\gamma_{\flat}\in A_{\Gamma}$ which have some features aligned with Theorem 2.9, and show that their product $\gamma_{\natural\sharp\flat}=\gamma_{\natural}\gamma_{\sharp}\gamma_{\flat}\in A_{\Gamma}$ is indeed a WPD contracting element. It may be useful for understanding the construction of $\gamma_{\natural\sharp\flat}$ to look at Section 5, which treats a concrete example.
@@
(4.1) $&V(Q(\Gamma,\Gamma_{0}))=\{V_{i}\mid 1\leq i\leq k\},\\ &E(Q(\Gamma,\Gamma_{0}))=\left\{(V_{i},V_{j})\middle|\begin{array}[]{l}1\leq i,\ j\leq k,\ i\neq j,\\ (\Gamma_{i})^{t}\text{ and }(\Gamma_{j})^{t}\text{ are connected in }(\Gamma_{0}\ast\Gamma_{i}\ast\Gamma_{j})^{t}\end{array}\right\}.\end{aligned}$
@@
  • (II)
    there exists a path from a vertex of $\Gamma_{i}$ and a vertex of $\Gamma_{j}$ on $\Gamma_{0}\ast\Gamma_{i}\ast\Gamma_{j}$, consisting of edges with label greater than $2$.
@@
For each $i,j$ with $i<j$ and $(V_{i},V_{j})\in E(T)$, take a path on $\Gamma$;
(4.2) $p_{i,j}=(w_{i,j,0},w_{i,j,1},\ldots,w_{i,j,d_{i,j}})$
consisting of edges with label greater than $2$ such that $w_{i,j,0}\in V_{i}$, $w_{i,j,d_{i,j}}\in V_{j}$ and $w_{i,j,1},\ldots,w_{i,j,d_{i,j}-1}\in V_{0}$. Moreover, set $d_{j,i}:=d_{i,j}$ and a path on $\Gamma$;
(4.3) $p_{j,i}=(w_{j,i,0},w_{j,i,1},\ldots,w_{j,i,d_{j,i}}):=(w_{i,j,d_{i,j}},\ldots,w_{i,j,1},w_{i,j,0}).$
For any $i,j\in\{1,\ldots,k\}$ with $(V_{i},V_{j})\in E(T)$, set
(4.4) $s_{i,j}:=w_{i,j,0}\in V_{i},\ t_{i,j}:=w_{i,j,d_{i,j}}\in V_{j}$
and
(4.5) $\tau_{i,j}=w_{i,j,0}w_{i,j,1}\cdots w_{i,j,d_{i,j}}\in A_{\Gamma}.$
Then we have $s_{j,i}=t_{i,j}$ in $V_{j}$, $t_{j,i}=s_{i,j}$ in $V_{i}$ and $\tau_{j,i}=w_{i,j,d_{i,j}}\cdots w_{i,j,1}w_{i,j,0}\in A_{\Gamma}.$
When $k\geq 2$, we take a closed path
(4.6) $(V_{i_{1}},\dots,V_{i_{r}},V_{i_{r+1}})$
with $i_{1}=i_{r+1}=1$ on the spanning tree $T$ of $Q(\Gamma,\Gamma_{0})$ passing through every vertex at least once. Since $r$ is the length of the closed path on the tree, $r$ is an even number. If $k=1$, then $T$ has no edges, and we set $r=0$.
To construct $\gamma_{\natural}$ and $\gamma_{\flat}$, we start with the following lemma, which is almost the same as [29, Lemma 6.1]. See Section 2.6 for notations related to matrices.